OpenStudy (anonymous):

What is a simpler form of the following expressions? (7p – 8)(7p + 8) (1 point) 49p2 + 112p + 64 49p2 – 112p – 64 49p2 + 64 49p2 – 64

OpenStudy (anonymous):

the last one

OpenStudy (anonymous):

Remember FOIL, First, Outside, Inside, Last. \[(7p-8)(7p+8)\] First \[7p*7p = 49p^{2}\] Outside \[7p*8 = 56p\] Inside \[-8*7p = -56p\] Last \[-8*8 = -64\] Add all these up \[49p^{2}+56p-56p-64 = 49p^{2}-64\] This also helps where if you have the form \[(Ax+b)(Ax-b) = A^{2}x^{2} - b^{2}\]